OpenStudy (anonymous):

Is anyone good at inscribed angles...I need homework help and I dont understand them at all

ganeshie8 (ganeshie8):

yeah post the q

OpenStudy (anonymous):

Segment AC is the diameter of the given circle.

OpenStudy (anonymous):

 @ganeshie8

ganeshie8 (ganeshie8):

notice that D is the center of circle.

ganeshie8 (ganeshie8):

since n\(\angle B\) is a semicircle angle, its 90

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

now use the angle sum property of triangle :- \(m\angle1 + m\angle2 + m\angle B = 180\) \(6x+2 + 16x + 90 = 180\)

ganeshie8 (ganeshie8):

simplify

ganeshie8 (ganeshie8):

solve x

OpenStudy (anonymous):

22x+92=180 -92 -92 22x=88 x=4 like that or am I way off

ganeshie8 (ganeshie8):

perfect !

OpenStudy (anonymous):

okay awesome!

ganeshie8 (ganeshie8):

so thats the x, but we need to find \(m\angle 1 \)

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

\(m\angle1 = 6x+2\)

ganeshie8 (ganeshie8):

plugin x value there

OpenStudy (anonymous):

okay so 6(4)+2 26?

ganeshie8 (ganeshie8):

thats right !

ganeshie8 (ganeshie8):

good work :)

OpenStudy (anonymous):

thanks!

ganeshie8 (ganeshie8):

np :)